x/2+y/4=6, x/5-y/2=0
1
Expert's answer
2014-07-21T13:53:01-0400
From the second equation of the system we conclude x/5=y/2, multiply by 2 and get y=2x/5. Next plug the last expression for y into the first equation of the system, so, x/2+(2x/5)/4=6, x/2+(x/5)/2=6, x/2+x/10=6. Multiply the last equation by 10 and obtain 5x+x=60, i.e., 6x=60, hence x=10. Therefore, y=2x/5, y=2*10/5, y=20/5, y=4.
Answer: x=10, y=4.